# Unreal Robotics Lab: A High-Fidelity Robotics Simulator with Advanced Physics and Rendering

### v1

@@ III-E Replay
Our system features an automated recording and replay mechanism for sensor and body pose data, enabling efficient data collection and experimentation. It uses custom structs to track only essential information –pose data, sensor readings, and timestamps– storing themin a queue for minimal overhead and precise sequence reproduction. Data streams are managed through separate JSON files, all of which can be dynamically saved and loaded at runtime. By publishing replayed data on the same ROS topics as live data, the system integrates seamlessly with existing pipelines. This allows researchers to collect data once and replay it in visually modified environments, supporting domain adaptation, robustness testing, and dataset augmentation for machine learning and benchmarking.
III-F Extended Features
Our system leverages Unreal Engine’s advanced rendering and simulation capabilities, surpassing traditional robotics simulators like MuJoCo [3], Gazebo [10], and Isaac Sim [5]. Key features include real-time global illumination (Lumen), high-detail mesh rendering (Nanite), and physically based reflections, which enhance perception-based robotics research. Procedural animations, NavMesh navigation, and built-in environmental tools enable efficient creation of dynamic simulations, while the Niagara particle system facilitates testing under adverse visual conditions (e.g., smoke, fog). Additionally, packaged standalone deployments improve reproducibility and scalability, addressing limitations of conventional simulators.
@@ III-G Example Deployment of Robots, Controllers, and Visual Planners
The system provides a unified framework where robots, controllers, and planners operate as in real-world deployments. All components communicate through ROS, ensuring seamless integration with external systems.
@@ III-G2 Controllers
The system includes joint controllers based on position and torque, as well as more advanced controllers such as Walk-These-Ways [26] and VBC [27]. UAVs use a PID-based controller, whereas manipulators rely on inverse kinematics (IK) methods such as Mink [28]. Since controllers operate within the ROS ecosystem, new control strategies can be easily integrated.
@@ III-G3 Planners
Motion planning is handled using learning-based and optimization-based planners such as ViNT [29], GNM [30], and NoMAD [31]. The ROS-based architecture of the framework ensures compatibility with additional external planners.
@@ III-G4 Visual SLAM
The system includes visual SLAM techniques such as OrbSLAM2 [32], OrbSLAM3 [33], and MASt3R-SLAM [34], enabling real-time localization and mapping. These methods integrate directly into the framework, enabling realistic perception and mapping in simulated environments.
@@ IV-A Simulation and Real-World Comparision
To evaluate the alignment between simulated and real-world visual inputs for image encoders, we employ Grad-CAM, EigenCAM, cosine similarity and KL Divergence. These methods assess both spatial attention and feature distribution similarity.
Grad-CAM computes the gradient of the class score $y^{c}$ with respect to feature map activations:
@@
Lower values indicate closer alignment.
To compare models trained on real-world data, we use CLIP [35] and the EfficientNet-B0 [36] visual encoder used in ViNT [29]. We first capture an image from the real world and recreate it within our simulator. Then both images are processed using Grad-CAM amd Eigen-Cam to generate attention heatmaps, as shown in Figure 4.
In both cases, the models primarily attend to the traffic cone in the center of the hallway as well as the end of the corridor, suggesting that the most salient features are preserved between the real and simulated environments. This alignment indicates that the models are focusing on the same high-level structures regardless of domain, reinforcing the realism of the simulated scene in terms of feature representation. The similarity in attention patterns also suggests that key objects remain distinguishable across domains, which is critical for transfer learning and sim-to-real applications.
To further quantify similarity, we compute KL divergence (0.4679) and cosine similarity (0.7157) between real and simulated feature distributions. Higher cosine similarity and lower KL divergence suggest closer alignment. While we did not aim to replicate real-world scenes with identical assets, opting instead for a simulation that remains visually representative of real conditions. Although the assets in our simulation differ from those in the real environment, our findings, together with GradCAM and EigenCAM comparisons, indicate that the simulation is visually representative enough for our evaluation purposes.
@@ IV-B Visual Navigation and SLAM Benchmarking
we benchmark standard methods to demonstrate simulator realism. We evaluated visual navigation (Experiment 1) with GNM [30] and ViNT [29], and visual SLAM (Experiment 2) with OrbSLAM2 [32], OrbSLAM3 [33], and MASt3R-SLAM [34]. Each method was run 5 times per condition, averaging results across trials. Both experiments used the Unitree Go1 quadruped with its onboard camera and the Walk-These-Ways low-level controller for realistic deployment.
The benchmarks were conducted in two distinct environments—warehouse and residential house—under three adversity levels: none, minor and major, classified by human judgment. UE’s Niagara system simulated environmental effects (smoke, rain, pollution, snow), while dynamic lighting and NPC movements were added for major adversity. Figure LABEL:fig:example_photo_realistic_render shows example environments and Figure 5 shows examples of adversity levels.
@@ IV-B1 Experiment 1: Visual Navigation
For visual navigation benchmarking, the robot was initialized at identical starting locations across environments and adversity levels. Each method had a fixed duration to reach the goal using a pre-built image-based topological map. The warehouse and residential house were used for this experiment. Methods were not fine-tuned, relying on a simplified topological map approach from the authors’ original repository [31], which may limit performance [37].
Success Weighted by Collision (SC). Following [38], we adopt the SC metric:
@@
TABLE I: Performance of different Visual navigation methods under various environments and varying adverse effects.
Table I presents the results, including Weighted SC to factor in collision penalties. State-of-the-art methods perform well under normal conditions, with ViNT generally outperforming GNM. Under minor adversity, GNM maintains a higher success rate but takes longer, while ViNT completes faster when successful. However, performance declines as environmental adversity increases (e.g., obstacles, smoke, rain), with both methods experiencing more collisions, longer times to goal, and lower SC. In severe adversity, neither method succeeds in the Warehouse, leading to zero SC. This suggests that these models struggle in extreme conditions due to limited training data, highlighting the need for training in diverse high-adversity conditions, which can be generated more effectively in simulation. The ability to create controlled but realistic adverse scenarios in a simulated environment provides a path to improve robustness in real-world deployments.
@@ IV-B2 Experiment 2: Visual SLAM
note that since monocular SLAM estimates trajectories with arbitrary scale, Umeyama’s method [39] is applied for scale and pose correction before computing coverage. Poor tracking or alignment can result in coverage exceeding 100%.
@@
Scaled $\text{Scaled ATE}=\frac{\text{ATE}}{\text{Coverage}}$ (8)
@@
TABLE II: Performance of Different VSLAM Methods under Varying Visual Effects.
From Table II and Fig. 6, Visual SLAM performance degrades with increasing visual adversity, highlighting sensitivity to challenging conditions. OrbSLAM2 struggles the most, with sharp declines in coverage and increased ATE, particularly in the warehouse environment. OrbSLAM3 performs well in clean conditions but deteriorates significantly under adversity, suggesting reduced robustness to occlusions, lighting changes, and dynamic elements. MASt3R-SLAM maintains the lowest ATE and highest coverage across conditions, demonstrating superior resilience to visual disturbances. Under severe visual adversity, all methods experience increased ATE and reduced coverage, highlighting the inherent challenges of maintaining reliable SLAM performance in dynamic or visually degraded environments. These replay-based evaluations provide a controlled, repeatable framework for assessing SLAM robustness across different levels of adversity, reinforcing the need for training and evaluation in diverse, high-adversity conditions to enhance real-world deployment reliability.
@@ V CONCLUSION
This paper presents a high-fidelity robotics simulation framework that integrates advanced photorealistic rendering with precise physics modeling to support vision-based robotics research and benchmarking. By leveraging Unreal Engine’s rendering capabilities and MuJoCo’s accurate physics, our system enables evaluation of robotic navigation and SLAM methods under diverse and adverse environmental conditions. Experimental results confirm that our simulator provides a robust testbed for evaluating real-world robustness in controlled, repeatable scenarios.
Despite its strengths, the system has certain limitations, including the absence of direct reinforcement learning (RL) support, lack of deformable terrain simulation, limited scene generation capabilities, and a focus on general robotics rather than autonomous vehicles or high-precision drone simulations.
Future Work. To address these limitations and further enhance our simulator, we plan to further enhance the MuJoCo integration, allowing the simulator to communicate with a standalone MuJoCo process and run RL environments (MuJoCo Playground [40]), generate synthetic datasets to enhance vision-based model training, expand automated scene generation, and enable imitation learning via VR-based teleoperation.
Overall, our platform serves as a robust foundation for benchmarking and data generation, bridging the gap between simulation and real-world robotics deployment.

### v2

@@ III-F Replay System
Our system has an automated recording and replay mechanism for sensor and body pose data, allowing efficient data collection and experimentation. It uses custom structs to track only essential information, pose data, sensor readings, and timestamps, storing them in a queue for minimal overhead and precise sequence reproduction. Data streams are managed through separate JSON files that can be dynamically saved and loaded at runtime. By publishing replayed data on the same ROS topics as live data, the system integrates seamlessly with existing pipelines. This allows researchers to collect data once and replay them in visually modified environments, supporting domain adaptation, robustness testing, and dataset enhancement for machine learning and benchmarking.
@@ III-G Example Deployment of Robots, Controllers, and Visual Planners
The system provides a unified framework in which robots, controllers, and planners operate as in real-world deployments. All components communicate through ROS, ensuring seamless integration with external systems.
@@ III-G2 Controllers
The system includes joint controllers based on position and torque, as well as more advanced controllers such as Walk-These-Ways [margolis2023walk] and VBC [liu2024visual]. UAVs use a PID-based controller, whereas manipulators rely on inverse kinematics (IK) methods such as Mink [Zakka_Mink_Python_inverse_2024]. Since controllers operate within the ROS ecosystem, new control strategies can be easily integrated.
@@ III-G3 Planners
Motion planning is handled using learning-based and optimization-based planners such as ViNT [shah2023vint], GNM [shah2023gnm], and NoMAD [sridhar2024nomad]. The ROS-based architecture of the framework ensures compatibility with additional external planners.
@@ III-G4 Visual SLAM
The system includes visual SLAM techniques such as OrbSLAM2 [murORB2], OrbSLAM3 [campos2021orb], and MASt3R-SLAM [murai2024mast3r], enabling real-time localization and mapping. These methods integrate directly into the framework, enabling realistic perception and mapping in simulated environments.
@@ IV-A Simulation and Real-World Comparision
To evaluate the alignment between simulated and real-world visual inputs for image encoders, we employ Grad-CAM, EigenCAM, cosine similarity and KL Divergence. These methods assess both spatial attention and feature distribution similarity. Grad-CAM computes the gradient of the class score $y^{c}$ with respect to feature map activations:
@@
Lower values indicate closer alignment. To compare models trained on real-world data, we use CLIP [radford2021learning] and the EfficientNet-B0 [tan2019efficientnet] visual encoder used in ViNT [shah2023vint]. We first capture an image from the real world and recreate it within our simulator. Then both images are processed using Grad-CAM amd Eigen-Cam to generate attention heatmaps, as shown in Figure 3. In both cases, the models primarily attend to the traffic cone in the center of the hallway as well as the end of the corridor, suggesting that the most salient features are preserved between the real and simulated environments. This alignment indicates that the models are focusing on the same high-level structures regardless of domain, reinforcing the realism of the simulated scene in terms of feature representation. The similarity in attention patterns also suggests that key objects remain distinguishable across domains, which is critical for transfer learning and sim-to-real applications. We further quantified the visual similarity between our hand-crafted simulations and real-world scenes by computing the KL divergence and cosine similarity for three scenes. These simulations were not built as photo-realistic replicas but as a proof of concept that our system can generate scenes visually representative of real conditions. As detailed in Table I, the Hallway scene, for example, yielded a KL divergence of 0.4679 and a cosine similarity of 0.7157, and an average of 0.4739 and 0.7094 in all scenes. These findings, along with our GradCAM and EigenCAM comparisons, confirm that our simulation is visually realistic enough for our evaluation purposes.
Metric Office Hallway Flat Hallway Bedroom Avg
KL divergence 0.4679 0.4435 0.5103 0.4739
Cosine similarity 0.7157 0.7231 0.6895 0.7094
TABLE I: Image comparison metrics between real-world scenes and simulated recreations.
@@ IV-B Visual Navigation and SLAM Benchmarking
We benchmark standard methods to demonstrate simulator realism. We evaluated visual navigation (Experiment 1) with GNM [shah2023gnm] and ViNT [shah2023vint], and visual SLAM (Experiment 2) with ORB-SLAM2 [murORB2], ORB-SLAM3 [campos2021orb], and MASt3R-SLAM [murai2024mast3r]. Each method was run 5 times per condition, averaging results across trials. Both experiments used the Unitree Go1 quadruped with its onboard camera and the Walk-These-Ways low-level controller for realistic deployment. The benchmarks were conducted in two distinct environments—warehouse and residential house—under three adversity levels: none, minor and major, classified by human judgment. UE’s Niagara system simulated environmental effects (smoke, rain, pollution, snow), while dynamic lighting and NPC movements were added for major adversity. Figure LABEL:fig:example_photo_realistic_render shows example environments and Figure 4 shows examples of adversity levels.
@@ IV-B1 Experiment 1: Visual Navigation
For visual navigation benchmarking, the robot was initialized at identical starting locations across environments and adversity levels. Each method had a fixed duration to reach the goal using a pre-built image-based topological map. The warehouse and residential house were used for this experiment. Methods were not fine-tuned, relying on a simplified topological map approach from the authors’ original repository [sridhar2024nomad], which may limit performance [visualnav-transformer-issue5].
Success Weighted by Collision (SC). Following [eftekhar2024one], we adopt the SC metric:
@@
TABLE II: Performance of different Visual navigation methods under various environments and varying adverse effects.
Table II presents the results, including Weighted SC to factor in collision penalties. State-of-the-art methods perform well under normal conditions, with ViNT generally outperforming GNM. Under minor adversity, GNM maintains a higher success rate but takes longer, while ViNT completes faster when successful. However, performance declines as environmental adversity increases (e.g., obstacles, smoke, rain), with both methods experiencing more collisions, longer times to goal, and lower SC. In severe adversity, neither method succeeds in the Warehouse, leading to zero SC. These failures are likely attributable to the dense, volumetric smoke simulated by the Niagara system, which caused persistent occlusion and aliasing of key visual features, preventing the vision-based planners from identifying a valid path. This suggests that these models struggle in extreme conditions due to limited training data, highlighting the need for training in diverse high-adversity conditions, which can be generated more effectively in simulation. The ability to create controlled but realistic adverse scenarios in a simulated environment provides a path to improve robustness in real-world deployments.
@@ IV-B2 Experiment 2: Visual SLAM
Note that since monocular SLAM estimates trajectories with arbitrary scale, Umeyama’s method [umeyama1991least] is applied for scale and pose correction before computing coverage. Poor tracking or alignment can result in coverage exceeding 100%.
@@
$\text{Scaled ATE}=\frac{\text{ATE}}{\text{Coverage}}$ (8)
@@
TABLE III: Performance of Different VSLAM Methods under Varying Visual Effects.
From Table III and Fig. 5, Visual SLAM performance degrades with increasing visual adversity, highlighting sensitivity to challenging conditions. ORB-SLAM2 struggles the most, with sharp declines in coverage and increased ATE, particularly in the warehouse environment. ORB-SLAM3 performs well in clean conditions but deteriorates significantly under adversity, suggesting reduced robustness to occlusions, lighting changes, and dynamic elements. MASt3R-SLAM maintains the lowest ATE and highest coverage across conditions, demonstrating superior resilience to visual disturbances. Under severe visual adversity, all methods experience increased ATE and reduced coverage, highlighting the inherent challenges of maintaining reliable SLAM performance in dynamic or visually degraded environments. These replay-based evaluations provide a controlled, repeatable framework for assessing SLAM robustness across different levels of adversity, reinforcing the need for training and evaluation in diverse, high-adversity conditions to enhance real-world deployment reliability.
IV-B3 Experiment 3: Real-world
Visual SLAM algorithms were evaluated on realworld tunnel dataset [kubelka2024we] under two conditions: clear and smoky, where a section of the tunnel was obscured by smoke generated from a smoke machine. Examples of these conditions are shown in Fig. 6. The percentage of tracked frames under these conditions was measured for various visual SLAM algorithms, results can be seen in Tab. IV, where it can be seen that the tracking performance of all SLAM algorithm degrades significantly under smoky conditions, consistent with the findings from our simulated experiments.
Env Adver. Frames tracked $\%$$\uparrow$
OrbSLAM2
Mono
OrbSLAM3
Mono
Mast3R
SLAM
Mine Clean 7.83 59.6 100
Smoky 0.00 38.5 43.6
TABLE IV: Performance of Different VSLAM Methods under Varying Visual Effects in the Real-World.
Figure 6: Examples of Realworld Tunnel data [kubelka2024we] used for evaluating VSLAM algorithm under adverse visual conditions.
IV-B4 Discussion
Our experiments collectively demonstrate that current state-of-the-art visual navigation and SLAM algorithms, while proficient in ideal conditions, remain brittle to the types of realistic, adverse visual phenomena that are produced by our framework and that are encountered in the realworld. The observed performance degradation highlights a critical gap in current training and testing paradigms. Futhermore, because adverse real-world datasets are scarce, there is a reinforced need for high-fidelity simulators such as URL, which can generate controlled, repeatable, and diverse high-adversity scenarios to advance the development of more robust robotic perception systems.
@@ V CONCLUSION
This paper presents a high-fidelity robotics simulation framework that integrates advanced photorealistic rendering with precise physics modeling to support vision-based robotics research and benchmarking. Using Unreal Engine’s rendering capabilities and MuJoCo’s accurate physics, our system enables evaluation of robotic navigation and SLAM methods under diverse and adverse environmental conditions. Experimental results confirm that our simulator provides a robust testbed for evaluating real-world robustness in controlled, repeatable scenarios. Despite its strengths, the system has certain limitations, including the absence of direct reinforcement learning (RL) support, lack of deformable terrain simulation, limited scene generation capabilities, and a focus on general robotics rather than autonomous vehicles or high-precision drone simulations. To address these limitations and further enhance our simulator, we plan to further enhance the MuJoCo integration, allowing the simulator to communicate with a standalone MuJoCo process and run RL environments (MuJoCo Playground [zakka2025mujocoplayground]), generate synthetic datasets to enhance vision-based model training, expand automated scene generation, and enable imitation learning via VR-based teleoperation. Overall, our platform serves as a robust foundation for benchmarking and data generation, bridging the gap between simulation and real-world robotics deployment.